# SparseWorld-TC: Trajectory-Conditioned Sparse Occupancy World Model

### v2

@@ 4 Experiments
DOME [dome] Occ GT 30.10 21.35 17.36 14.86 12.61 11.03 10.00 9.34 15.83
COME [come] Occ GT 33.78 24.57 21.35 18.25 15.84 13.85 12.99 11.96 19.07
@@
DOME [dome] Occ GT 39.04 31.20 27.14 24.73 22.32 20.28 19.05 17.97 25.21
COME [come] Occ GT 44.20 36.25 32.86 30.03 26.93 24.70 23.30 21.44 29.96
@@ 4.1 Experimental Setup
Dataset and Metrics. Experiments are conducted on the widely adopted Occ3D-nuScenes benchmark [occ3d]. Our implementation strictly follows the experimental settings established in previous works [dome, occworld, dfit-occworld, come, occ-llm, i2world]. For evaluation, we adopt the standard geometric Intersection over Union (IoU) and semantic mean Intersection over Union (mIoU) metrics.
Implementation Details. As detailed in the methodology section, our architecture includes several configurable components. We pick past frames within the past 2 seconds, following recent models [dome, occ-llm, dfit-occworld, occworld, i2world, come]. The future prediction horizon $T$ is set to 3s and 8s, aligning with established benchmarks [occworld, dome]. To accommodate varying computational constraints, we configure the number of anchors per frame as $N=600$, with each anchor comprising $M=128$ points. This configuration, designated as SparseWorld-TC-Small, adopts a ResNet-50 backbone [resnet]. Our SparseWorld-TC-Large variant uses $N=4800$ and $M=16$, delivering improved performance at increased computational cost. Furthermore, leveraging the representational power of DINOv3 [dinov3], we introduce SparseWorld-TC-Small* and SparseWorld-TC-Large*, which adopts a DINOv3-Base image backbone while retaining the same configurations as “-Small” and “-Large” versions. All models are trained for 70 epochs on 8 NVIDIA H20 GPUs with a total batch size of 8, using the AdamW optimizer [adamw] with an initial learning rate of $2\times 10^{-4}$ and a cosine annealing learning rate schedule [lr].
@@ 4.2 Main Results
4D Occupancy Forecasting. We evaluate our proposed method following established evaluation protocols [dome, occworld, dfit-occworld, i2world], as summarized in Table 1. While some VAE-based methods report performance using ground-truth occupancy as input, a requirement sometimes unavailable in real-world scenarios, we include these results for a comprehensive comparison. Since our approach directly consumes camera sensor input rather than precomputed occupancy grids, we primarily compare against camera-based versions of prior methods. Note that OccWorld-D uses predictions from TPVFormer [tpvformer]; “-STC” and “-F” indicate models incorporating occupancy results from STCOcc [stcocc] and FB-Occ [fbocc], respectively.
Our SparseWorld-TC-Large achieves significant improvements, outperforming the previous state-of-the-art by 18.7% in mIoU (26.42 vs. 22.26) and 11.7% in IoU (49.21 vs. 44.07). Due to the flexible and variable number of queries, our model exhibits strong scalability. SparseWorld-TC-Small achieves 25.60 mIoU and 49.02 IoU, while being approximately twice as fast as SparseWorld-TC-Large and I2-World [i2world], thus striking an effective balance between performance and efficiency. The strong performance of the DINO version of our method further demonstrates the scalability in leveraging large-scale foundation models.
Moreover, although Table 1 shows that methods using ground-truth occupancy generally achieve better performance than those processing raw sensor inputs, our camera-based model remains competitive across several metrics. Specifically, SparseWorld-TC-Large* achieves a higher average mIoU (29.89 vs. 27.10) than DOME [dome] and a higher average IoU (53.52 vs. 49.80) compared to I2-World [i2world], which can be attributed to its reduced decay in long-term occupancy forecasting.
@@ 4.3 Ablation Study
Effect of Trajectory. Our model supports the conditioning of arbitrary future trajectories. Through ablation studies, we examine how such trajectory guidance influences model behavior and output quality. As summarized in Table 3, we evaluate performance in three settings based on the SparseWorld-TC-Small model: without trajectory input, with trajectories predicted by BEVPlanner [bev-planner], and with ground-truth trajectories. The results demonstrate that our model remains robust even when provided with predicted trajectories, while performance consistently improves with more accurate trajectory estimates.
@@ 4.4 Visualizations & Discussions
From Occupancy to Sensors. Building upon the occupancy world model, our proposed architecture is further capable of directly predicting future sensor observations, adopting an extension inspired by [infinicube, scube, drivingforward]. This capability is realized by incorporating additional MLPs into the decoding network illustrated in Figure 2. All of these MLPs collectively generate 3D Gaussian parameters, center offsets, and classification logits in a single feed-forward pass. For a comprehensive exposition of our methodology and additional experimental results, please refer to the supplementary document.
@@ 5 Conclusion
[…]
@@ A.1 Ray-level mIoU
SparseOcc [sparseocc] proposes RayIoU (Ray-level mIoU) to solve the inconsistency penalty along the depth axis raised in traditional voxel-level mIoU criteria. We evaluate our SparseWorld-TC-Large* model with this ray-level metric and report the results in Table 5.
Table 5: RayIoU scores $\%$] of 4D occupancy forecasting performance on the Occ3D-nuScenes [occ3d] benchmark.
@@ A.2 Ablation study on trajectory embedding
A.3 Per-class performance
@@
Table 7: Per-class mIoU $\%$] performance of 4D occupancy forecasting on Occ3D-nuScenes [occ3d].
@@
Table 8: Per-class RayIoU $\%$] performance of 4D occupancy forecasting on Occ3D-nuScenes [occ3d].
@@
B Additional Qualitative Experiments
B.1 Feedforward Gaussian
Inspired by advances in feedforward Gaussian methods [scube, infinicube, drivingforward], we extend the original model with additional MLPs to decode Gaussian parameters from latent features. Then we utilize the differential Gaussian rasterization proposed in 3DGS [3dgaussian] to render the predicted front-view image and calculate L1 loss with the GT image.
GT Pred.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Gaussian splatting reconstruction during training.
GT Pred.
1s
Refer to caption Refer to caption
2s
Refer to caption Refer to caption
3s
Refer to caption Refer to caption
Figure 8: Future observation forecasting on validation set.
B.2 Additional Trajectory-conditioned Prediction
More visualizations are shown below.

### v3

@@ 4 Experiments
DOME [8] Occ GT 30.10 21.35 17.36 14.86 12.61 11.03 10.00 9.34 15.83
COME [34] Occ GT 33.78 24.57 21.35 18.25 15.84 13.85 12.99 11.96 19.07
@@
DOME [8] Occ GT 39.04 31.20 27.14 24.73 22.32 20.28 19.05 17.97 25.21
COME [34] Occ GT 44.20 36.25 32.86 30.03 26.93 24.70 23.30 21.44 29.96
@@ 4.1 Experimental Setup
Dataset and Metrics. Experiments are conducted on the widely adopted Occ3D-nuScenes benchmark [38]. Our implementation strictly follows the experimental settings established in previous works [8, 55, 53, 34, 49, 25]. For evaluation, we adopt the standard geometric Intersection over Union (IoU) and semantic mean Intersection over Union (mIoU) metrics.
Implementation Details. As detailed in the methodology section, our architecture includes several configurable components. We pick past frames within the past 2 seconds, following recent models [8, 49, 53, 55, 25, 34]. The future prediction horizon $T$ is set to 3s and 8s, aligning with established benchmarks [55, 8]. To accommodate varying computational constraints, we configure the number of anchors per frame as $N=600$, with each anchor comprising $M=128$ points. This configuration, designated as SparseWorld-TC-Small, adopts a ResNet-50 backbone [10]. Our SparseWorld-TC-Large variant uses $N=4800$ and $M=16$, delivering improved performance at increased computational cost. Furthermore, leveraging the representational power of DINOv3 [35], we introduce SparseWorld-TC-Small* and SparseWorld-TC-Large*, which adopts a DINOv3-Base image backbone while retaining the same configurations as “-Small” and “-Large” versions. All models are trained for 70 epochs on 8 NVIDIA H20 GPUs with a total batch size of 8, using the AdamW optimizer [31] with an initial learning rate of $2\times 10^{-4}$ and a cosine annealing learning rate schedule [30].
@@ 4.2 Main Results
4D Occupancy Forecasting. We evaluate our proposed method following established evaluation protocols [8, 55, 53, 25], as summarized in Table 1. While some VAE-based methods report performance using ground-truth occupancy as input, a requirement sometimes unavailable in real-world scenarios, we include these results for a comprehensive comparison. Since our approach directly consumes camera sensor input rather than precomputed occupancy grids, we primarily compare against camera-based versions of prior methods. Note that OccWorld-D uses predictions from TPVFormer [11]; “-STC” and “-F” indicate models incorporating occupancy results from STCOcc [24] and FB-Occ [21], respectively.
Our SparseWorld-TC-Large achieves significant improvements, outperforming the previous state-of-the-art by 18.7% in mIoU (26.42 vs. 22.26) and 11.7% in IoU (49.21 vs. 44.07). Due to the flexible and variable number of queries, our model exhibits strong scalability. SparseWorld-TC-Small achieves 25.60 mIoU and 49.02 IoU, while being approximately twice as fast as SparseWorld-TC-Large and I2-World [25], thus striking an effective balance between performance and efficiency. The strong performance of the DINO version of our method further demonstrates the scalability in leveraging large-scale foundation models.
Moreover, although Table 1 shows that methods using ground-truth occupancy generally achieve better performance than those processing raw sensor inputs, our camera-based model remains competitive across several metrics. Specifically, SparseWorld-TC-Large* achieves a higher average mIoU (29.89 vs. 27.10) than DOME [8] and a higher average IoU (53.52 vs. 49.80) compared to I2-World [25], which can be attributed to its reduced decay in long-term occupancy forecasting.
@@ 4.3 Ablation Study
Effect of Trajectory. Our model supports the conditioning of arbitrary future trajectories. Through ablation studies, we examine how such trajectory guidance influences model behavior and output quality. As summarized in Table 3, we evaluate performance in three settings based on the SparseWorld-TC-Small model: without trajectory input, with trajectories predicted by BEVPlanner [23], and with ground-truth trajectories. The results demonstrate that our model remains robust even when provided with predicted trajectories, while performance consistently improves with more accurate trajectory estimates.
@@ 4.4 Visualizations & Discussions
From Occupancy to Sensors. Building upon the occupancy world model, our proposed architecture is further capable of directly predicting future sensor observations, adopting an extension inspired by [32, 33, 37]. This capability is realized by incorporating additional MLPs into the decoding network illustrated in Figure 2. All of these MLPs collectively generate 3D Gaussian parameters, center offsets, and classification logits in a single feed-forward pass. For a comprehensive exposition of our methodology and additional experimental results, please refer to the supplementary document.
@@ 5 Conclusion
Acknowledgments
This paper is supported by the National Natural Science Foundation of China under Grants 62233013.
@@ A.1 Ray-level mIoU
SparseOcc proposes RayIoU (Ray-level mIoU) to solve the inconsistency penalty along the depth axis raised in traditional voxel-level mIoU criteria. We evaluate our SparseWorld-TC-Large* model with this ray-level metric and report the results in Table 5.
Table 5: RayIoU scores $\%$] of 4D occupancy forecasting performance on the Occ3D-nuScenes [38] benchmark.
@@ A.2 Ablation Study on Trajectory Embedding
A.3 Ablations of More Model Settings
In Table 7, “w/o Deformable Attn” replaces deformable attention with a simpler multi-view feature aggregation, while “w/o Temporal Attn” replaces temporal attention with frame-wise processing. We further ablate the effect of increasing number of anchors and points per anchor in Table 8. Roughly, more anchors and points per anchor both improve the performance.
Table 7: Ablation on Removing Temporal and Deformable Attention. Avg. denotes average of mIoU or IoU in 1s, 2s, and 3s.
Modules Avg.mIoU ($\%$) $\uparrow$ Avg.IoU ($\%$) $\uparrow$
Ours-Small 25.60 49.02
w/o Temporal Attn. 25.07 47.53
w/o Deformable Attn. 23.89 46.55
A.4 Hardware Cost versus Model Size and Performance
As shown in Table 8, the + denotes increasing the number of anchors (N) or points per anchor (M). We compare on backbones with both ResNet and DINOv3(*).
Table 8: GPU Memory Usage and Performance of Evaluation.
Model N M Memory (GB) mIoU ($\%$)
Ours-Small 600 128 5.4 25.60
Ours-Large 4800 16 7.7 26.42
Ours-Large+ 4800 64 9.2 27.25
Ours-Small* 600 128 7.1 27.71
Ours-Large* 4800 16 14.4 29.89
Ours-Large*+ 9600 16 24.6 30.88
A.5 Per-class Performance
@@
A.6 Improvement on Small Targets
The performance on smaller objects, including motorcycles and pedestrians, improves with a larger number of anchor queries. Emphasizing small objects through loss re-weighting also enhances the model capability, as evidenced in Table 9.
Table 9: Performance Improvement for Small Objects
Model Avg. mIoU ($\%$) $\uparrow$
motorcycle pedestrian
Ours-Large* 13.64 6.86
Q=4800 $\rightarrow$ Q=9600 14.72 (+1.08) 8.29 (+1.43)
Class Reweighting 14.64 (+1.00) 8.63 (+1.83)
B Additional Qualitative Experiments
B.1 Feedforward Gaussian
Inspired by advances in feedforward Gaussian methods [33, 32, 37], we extend the original model with additional MLPs to decode Gaussian parameters from latent features. Then we utilize the differential Gaussian rasterization proposed in 3DGS [14] to render the predicted front-view image and calculate L1 loss with the GT image.
Currently our feedforward 3DGS implementation only supports the front-view camera, which supports both 256x704 and 128x352 resolutions. To alleviate the voids caused by query sparsity, we boost the number of anchors and points per anchor to 7200 and 32.
GT Pred.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Gaussian splatting reconstruction during training.
GT Pred.
1s
Refer to caption Refer to caption
2s
Refer to caption Refer to caption
3s
Refer to caption Refer to caption
Figure 8: Future observation forecasting on validation set.
B.2 Convergence Visualization
Figure 9 shows how the model converge from noise to the detailed occupancy.
Figure 9: Model Convergence Visualization.
B.3 Failure Case Analysis
Figure 10: Failure cases: (left) the car and barrier with heavy occlusion, (right) small objects appear in the future 6 seconds.
As shown in Fig. 10, we visualize typical failure cases.
B.4 Additional Trajectory-conditioned Prediction
We additionally visualize the occupancy prediction results under different trajectory conditions, as shown in Fig. 11. More visualizations are shown in Fig. 12 and Fig. 13.
Table 10: Per-class mIoU $\%$] performance of 4D occupancy forecasting on Occ3D-nuScenes [38].
@@ B.4 Additional Trajectory-conditioned Prediction
Table 11: Per-class RayIoU $\%$] performance of 4D occupancy forecasting on Occ3D-nuScenes [38].